jane travels 6 time faster than bill traveling in opposite direction they are 378 miles apart after 4.5 hours find their rates of travel
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

jane travels 6 time faster than bill traveling in opposite direction they are 378 miles apart after 4.5 hours find their rates of travel
------
Bill DATA:
rate = x mph ; time = 4.5 hrs ; distance = 4.5x miles
----
Jane DATA:
rate = 6x mph ; time = 4.5 hrs ; distance = 4.5*6x = 27x miles
-----
Equation:
dist + dist = 378 miles
4.5x + 27x = 378
31.5x = 378
-----
x = 12 mph (Bill's rate)
6x = 72 mph (Jane's rate)
--------
Cheers,
Stan H.
---------

Answer by Alan3354(69443)

Is it 6 times faster?
Or 6 times as fast?